40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 16 14:56:44 2021, gautam, Update, SUS, aLIGO Sat Amp installed, powered MC1_deadUL.pngIMG_9100.jpg
    Reply  Wed Feb 17 16:11:53 2021, gautam, Update, SUS, aLIGO Sat Amp installed, powered 
       Reply  Thu Feb 18 15:33:21 2021, gautam, Update, SUS, aLIGO Sat Amp installed, powered and commissioned 
          Reply  Thu Feb 18 18:05:04 2021, gautam, Update, SUS, aLIGO Sat Amp characterization OSEMnoise.pdfLEDdriveNoise.pdf
             Reply  Thu Feb 18 20:20:25 2021, Koji, Update, SUS, aLIGO Sat Amp characterization 
Message ID: 15809     Entry time: Tue Feb 16 14:56:44 2021     Reply to this: 15814
Author: gautam 
Type: Update 
Category: SUS 
Subject: aLIGO Sat Amp installed, powered 

[jordan, gautam]

  • Ran 60ft long cables from 1X4 to MC1/MC3 chamber flange, via overhead cable tray, and top of PSL enclosure for the last ~20ft. Note that it may be that the overhead cable trays cannot support the weight of the cables for 15 SOSs (total 30 shielded cables with 37 wires as twisted pairs) when we eventually add the optics for the BHD upgrade.
  • Installed aLIGO satellite amplifer in 1X4.
  • Tapped +/-20 V (which is the available voltage closest to the required +/-18V). For this, the Sorensens were powered down, and the actual taps were made from the fusable blocks powering the Trillium interface box. We made sure to leave an extra slot so that this kind of additional headache is not required for the next person doing such work.
  • Once installed, I plugged in the dummy suspension box and verified that the unit performs as expected. 
  • Some photos of the installation are here.

After this work, the IMC locked fine, the AS camera has the Michelson fringing, the fast CDS indicators are all green, and the seismometer BLRMS all look good - therefore, I claim no lasting damage was done as a direct result of today's work at 1X4. I will connect up the actual suspension at my leisure later today. Note that the MC1 glitches seem to have gone away, without me doing anything about it. Nevertheless, I think it's about time that we start testing the new hardware. 


Unrelated to this work: while I was testing some characteristics of the MC1 suspension (before we did any work in the VEA, you can see the timestamp in the ndscope), I noticed that the MC1 UL coil channel cannot actually be used to actuate on the optic. The coil driver Vmon channel demonstrates the appropriate response, which means that the problem is either with the Satellite box (it is just a feedthrough, so PCB trace damaged?) or with the OSEM itself (more likely IMO, will know more once I connect the new Satellite Amplifier up). I only show comparison for UL vs UR, but I checked that the other coils seem to be able to actuate the optic. This means we have been running for an indeterminate amount of time with only 3 face actuators on MC1, probably related to me having to do this work


Also unrelated to this work - while poking around at 1X5 rear, I noticed that the power connections to the existing Satellite Boxes are (understatedly) flaky, see connections to T1-T4 in Attachment #2..

Attachment 1: MC1_deadUL.png  202 kB  | Show | Hide all | Show all
Attachment 2: IMG_9100.jpg  3.046 MB  Uploaded Tue Feb 16 16:38:38 2021  | Hide | Hide all | Show all
IMG_9100.jpg
ELOG V3.1.3-